# Towards Atomic MIMO Receivers

### v2

@@ I Introduction
With the global deployment of commercial 5G services, cutting-edge research on 6G has been launched worldwide [1]. Meanwhile, the revolutionary quantum information technologies are evolving rapidly and deemed as a key enabler of the next-generation computing, communication, and sensing [2, 3, 4]. For example, ever since the invention of Schor’s algorithm, quantum computing has exhibited unparalleled abilities in quadratically or even exponentially accelerating the computing speed in solving classically intractable problems, ranging from factorization to optimization to artificial intelligence [2]. On the other hand, leveraging quantum mechanisms such as superposition and entanglement, quantum communication is capable of establishing unconditionally secure links to protect data privacy [3]. In parallel with communication and computing, quantum sensing is another promising branch of quantum information technologies [4]. It employs microscopic particles as “sensors” and capitalizes on their strong sensitivity to external disturbance, especially electromagnetic fields, to precisely measure some physical quantities. Notable examples of quantum sensors include atomic clocks and superconducting interferometers. As an exciting trend, quantum sensors are being applied to an increasingly broad range of practical use cases [4, 5, 6]. Relevant techniques promise to reshape the information society and our perception of the environment. Aligned with this end, our work advocates the use of quantum sensors to revolutionize next-generation wireless receivers.
@@ I-A Concept and Properties of Atomic Receiver
Benefiting from the advancements in quantum sensing, atomic receivers offer several advantages over conventional receivers for wireless communication and sensing. First, atomic receivers can achieve much higher detection-and-sensing accuracies as mentioned earlier [4]. For example, as demonstrated in [5], atomic receivers can achieve the sub-millimeter (sub-mmWave) level granularity in detecting object vibrations using WiFi sources as opposed to the mmWave level granularity attained by traditional receivers. The second advantage lies in low thermal noise. Conventional receivers require RF circuits to intercept radio waves and downconvert RF signals to the baseband, which comprise components such as RF antennas, analog filters, and mixers that generate significant thermal noise. In contrast, atomic receivers employ all-optical components for sensing and downconversion, and thus are immune to thermal noise [9]. The above two advantages make atomic receivers a promising technology for long-range communications [6]. Last but not the least, the flexible tunability of Rydberg atoms across an ultra-wide spectrum is a significant advantage [13]. Traditional antennas, with a fixed size typically half the wavelength of the carrier, are limited in the spectrum bandwidth they can listen to. Detecting multi-band signals requires an array of costly RF receivers, each of which is matched to a specific band [14]. In contrast, the numerous energy levels in Rydberg atoms allow an atomic receiver to respond to a broad range of frequencies from 100 Megahertz (MHz) to 1 Terahertz (THz) by exciting electrons to different energy levels [15]. This capability enables the simultaneous measurement of multi-band signals over an extremely wide spectrum using a single atomic receiver, laying the foundation for constructing an extraordinarily cost-effective full-frequency communication platform. To materialize the above visions, this work incorporates atomic receivers into modern wireless communication systems.
@@ I-B State-of-the-art Atomic Receivers
Despite the rapid development of atomic receivers in the quantum-sensing domain, their application in wireless communication is still at a nascent stage. The state-of-the-art designs are suitable only for primitive communication scenarios and schemes, such as a point-to-point system or a single-input-single-output (SISO) air interface. Without advanced designs and advices, the full potential of atomic receivers cannot be unleashed, with the state-of-the-art achieving only a Mbps-scale channel capacity [20, 15, 22]. Communication techniques that are customized for atomic receivers are largely uncharted. This calls for research on atomic-level channel coding, massive multiple-input-multiple-output (MIMO), multi-user communications, integrated sensing and communication (ISAC), reconfigurable intelligent surface (RIS), and edge learning [23, 24, 25]. To contribute to this emerging field, this paper seeks the feasibility of integrating MIMO communications and atomic receivers. MIMO, a key 5G technology, features spatial multiplexing of parallel data links and enhancement of link reliability via diversity gain [23]. It is believed that the consideration of MIMO has the potential to overcome the capacity bottleneck of current atomic SISO receivers.
@@ I-C Contributions and Organization
  • Performance study: We employ two metrics, namely the normalized mean square error (NMSE) and the bit error ratio (BER), to validate the feasibility of atomic MIMO receivers as well as the effectiveness of the proposed algorithms. Targeting the NMSE metric, the Cramér-Rao lower bound (CRLB) is derived. Numerical results demonstrate that the proposed EM-GS algorithm consistently approaches the CRLB and outperforms the biased GS algorithm, especially in the low signal-to-noise ratio (SNR) regime. Regarding the BER performance, simulation results confirm that the biased GS and EM-GS algorithms achieve BERs close to those from an exhaustive search, while being much more computational efficient.
The remainder of this paper is organized as follows. The preliminaries of atomic SISO receivers are provided in Section II. The framework of atomic MIMO receiver and the corresponding detection algorithms are presented in Section III and Section IV, respectively. In Section V, we validate the proposed design by numerical results. Finally, conclusions are drawn in Section VI.
@@ II-C1 Quantum State of Rydberg Atom
${\mathbf{E}}(t)=\bm{\epsilon}\sqrt{P}\rho s\cos(\omega t+\varphi),$ (2)
@@ II-C2 State Evolution of Rydberg Atom
    $\hat{V}=q\sum_{d\in\{x,y,z\}}\hat{r}_{d}E_{d}(t)=\left(\begin{array}[]{cc}0&\xi(t)\\ \xi^{*}(t)&0\end{array}\right),$ (8)
@@
where $\delta\overset{\Delta}{=}\omega_{eg}-\omega$ denotes the detuning of the carrier frequency $\omega$ from the transition frequency $\omega_{eg}$. In (9), $\Omega_{R}\overset{\Delta}{=}\frac{|({\mathbf{r}}_{eg}^{T}\bm{\epsilon})q\sqrt{P}\rho s|}{\hbar}$ denotes the Rabi frequency, defined as the frequency at which the probability amplitudes of two atomic energy levels fluctuate in response to an oscillating electromagnetic field [29]. For ease of notation, we introduce $\Omega\overset{\Delta}{=}\sqrt{\Omega^{2}_{R}+\delta^{2}}$ to represent the effective Rabi frequency.
@@ II-E Implementation of Atomic Receiver
Despite the elegant mathematical representation of Rabi oscillation in (9), observing this phenomenon in practice is quite challenging due to various external disturbances that can destroy the coherence of a quantum state, such as spontaneous emission [29]. Fortunately, the data-bearing effective Rabi frequency is still measurable, as it can induce two other quantum phenomena, namely electromagnetically induced transparency (EIT) and Autler-Townes (AT) splitting [6, 8]. As shown in Fig. 2, the EIT-AT phenomenon involves a four-level system denoted by quantum states |$\ket{1}$, |$\ket{2}$, |$\ket{3}$, and |$\ket{4}$, where |$\ket{3}$ and |$\ket{4}$ represent the Rydberg states we studied before. Recall that when preparing a Rydberg state, two laser beams, the probe beam and coupling beam, are needed to excite the electron from |$\ket{1}$ to |$\ket{2}$ and then from |$\ket{2}$ to |$\ket{3}$. In practice, a PD is deployed to receive the probe beam and illustrate its spectrum for observing the Rabi frequency. When the incident electromagnetic wave is turned off and the probe beam is turned resonant with the transition |$\ket{1}\rightarrow\ket{2}$, the EIT phenomenon occurs such that the probe beam penetrates through the vapor cell nearly without absorption loss. What reflects on its spectrum is that one peak appears at the resonant frequency (see the upper sub-figure of Fig. 2). On the other hand, when the incident electromagnetic wave is turned on, the two Rydberg states |$\ket{3}$ and |$\ket{4}$ are coupled by the Rabi frequency. As a result, the AT splitting phenomenon is induced, which splits the peak of the spectrum of the probe beam into two as depicted in the lower sub-figure of Fig. 2. As physicists have demonstrated, the splitting interval $\Delta f$ of the two peaks is linearly proportional to the effective Rabi frequency [9]:
@@ III Overview of Atomic MIMO Receiver
In the preceding section, we introduce the principles of atomic SISO receiver. In this section, building on these principles, we propose the atomic MIMO receiver for multi-user communications. First, its architecture and principles are presented. Then, the problem of atomic-MIMO detection is formulated, which is solved in the next section.
III-A Architecture and Principles of Atomic MIMO Receiver
Figure 3: Schematic diagram of the atomic MIMO receiver. One atomic receiver employing $N$ atomic antennas is measuring signals from $K$-single antenna users together with one reference source, polluted by external noise.
Consider the multi-user communication system illustrated in Fig. 3, where $K$ single-antenna users communicate with one atomic MIMO receiver. Its architecture comprises $N$ vapor cells filled with Rydberg atoms and an equal number of PDs. Each vapor cell acts as an atomic antenna for measuring electromagnetic waves, while the PDs read out the measurement results (i.e., the splitting interval, $\Delta_{f}$, and the Rabi frequency, $\Omega_{R}$) and aggregate them to the digital processor for signal detection. The simultaneous users share the same carrier frequency $\omega$, while all atoms are triggered to the uniform Rydberg states $\hbar\omega_{g}$ and $\hbar\omega_{e}$ (e.g., $52D_{5/2}$ and $53P_{3/2}$) in response to the carrier frequency (e.g., $\omega=5$ GHz). As a result, the resonance case with $\omega=\omega_{eg}$ is created. The principles and operations of the receiver are described in the sequel.
Consider an arbitrary user, say the $k$-th user. Its phase modulated symbol is denoted as $s_{k}=|s_{k}|e^{i\gamma_{k}}\in\mathcal{S}$, where the quadrature amplitude modulation (QAM) is adopted and the set $\mathcal{S}$ corresponds to the QAM constellation. The average power of $s_{k}$ is normalized as $\mathsf{E}(\|s_{k}\|^{2})=1$. Then, the transmit RF signal is given by $x_{k}(t)=\sqrt{P_{k}}|s_{k}|\cos(\omega t+\gamma_{k})$. We adopt the practical multi-path wireless channel with $L$ paths. The simultaneous transmissions from all users result in the received electromagnetic wave at the $n$-th vapor cell:
@@ III-A Architecture and Principles of Atomic MIMO Receiver
$\color[rgb]{0,0,0}\xi_{n}(t){=}\sum_{k=1}^{K}\sum_{l=1}^{L}{{\mathbf{r}}}_{eg}^{T}\bm{\epsilon}_{nkl}q\sqrt{P_{k}}\rho_{nkl}|s_{k}|\cos(\omega t+\varphi_{nkl}+\gamma_{k}).$ (14)
@@
$\color[rgb]{0,0,0}\Omega_{R,n}=\left|\sum_{k=1}^{K}\sum_{l=1}^{L}\frac{1}{\hbar}({{\mathbf{r}}}_{eg}^{H}\bm{\epsilon}_{nkl})q\sqrt{P_{k}}\rho_{nkl}|s_{k}|e^{i(\varphi_{nkl}+\gamma_{k})}\right|.$ (16)
@@
The goal of atomic-MIMO receiver is to read the values of $\{\Omega_{R,n}\}$ and infer $\{s_{k}\}$ from $\{\Omega_{R,n}\}$. To this end, according to the principle of EIT-AT discussed in Section II-E, the Rabi frequency, $\Omega_{R,n}$, can be measured by the splitting interval, $\Delta f_{n}$, of the probe-beam spectrum since $\Omega_{R,n}=2\pi\frac{\lambda_{p}}{\lambda_{c}}\Delta f_{n}$. For ease of exposition, let the measurement of $\Omega_{R,n}$ be denoted as $z_{n}$, which is the equivalent received signal at the $n$-th antenna. To facilitate the detector design, a compact matrix form of $\{z_{n}\}$ is derived as follows. We define ${\mathbf{s}}=[s_{1},s_{2},\cdots,s_{K}]^{T}$ and ${\mathbf{a}}_{n}=[a_{n,1},a_{n,2},\cdots,a_{n,K}]^{T}$, where $a_{n,k}=\sum_{l=1}^{L}\frac{1}{\hbar}({{\mathbf{r}}}_{eg}^{T}\bm{\epsilon}_{n,k})q\sqrt{P_{k}}\rho_{nkl}e^{-i\varphi_{nkl}}$. It is straightforward to obtain that $z_{n}=\left|{\mathbf{a}}_{n}^{H}{\mathbf{s}}\right|$. We further define the received vector-form signal as ${\mathbf{z}}=[z_{1},z_{2},\cdots,z_{N}]^{T}$ and the observation matrix as ${\mathbf{A}}=[{\mathbf{a}}_{1},{\mathbf{a}}_{2},\cdots,{\mathbf{a}}_{N}]\in\mathbb{C}^{K\times N}$. It follows that
@@
where the operator $|{\mathbf{x}}|$ takes the magnitude of each element of the vector, ${\mathbf{x}}$. A crucial conclusion is drawn from (17) that the transmit symbols in ${\mathbf{s}}$ are embedded in the observed amplitude vector ${\mathbf{z}}=|{\mathbf{A}}^{H}{\mathbf{s}}|$, where the observation matrix ${\mathbf{A}}$ can be treated as an effective MIMO channel. As opposed to the linear model adopted in conventional MIMO communications, the atomic MIMO receiver features a non-linear model based on amplitude detection. This is equivalent to the well-known PR model in the realm of optical imaging [30].
Before delving into this PR problem, two factors, which are part of the proposed framework and affecting the receiver effectiveness, are described as follows.
  • Reference signal: To begin with, as shown in Fig. 3, a measurement scheme called holographic phase-sensing, where a LO sends a known reference signal, $s_{b}$, to the atomic receiver, is often employed to detect QAM symbols [19]. To be specific, as the phase information is erased in (17), the absolute phase $e^{i\angle{\mathbf{s}}}$ of ${\mathbf{s}}$ is undetectable. However, with the help of the known reference signal $s_{b}$, one can infer the absolute phase, $e^{i\angle{\mathbf{s}}}$, by estimating the phase difference between ${\mathbf{s}}$ and $s_{b}$, thereby enabling QAM detection. Let the reference signal have the same carrier frequency $\omega$222 Existing atomic SISO receivers also widely adopt the quantum heterodyne detection, where the frequency of reference signal slightly differs from the user signal. This article concentrates on the holographic-phase sensing measurement, while the corporation of heterodyne detection with atomic MIMO receiver is left for future works.. Then, following the same steps as (11) to (17), we can arrive at the modified received signal: ${\mathbf{z}}=|{\mathbf{A}}^{H}{\mathbf{s}}+{\mathbf{b}}|$, where ${\mathbf{b}}=[b_{1},b_{2},\cdots,b_{N}]^{T}$, $b_{n}=\frac{s_{b}}{\hbar}({{\mathbf{r}}}_{eg}^{H}\bm{\epsilon}_{b,n})q\sqrt{P_{b}}\rho_{b,n}e^{i\varphi_{b,n}}$, and $P_{b}$, $\rho_{b,n}$, $\varphi_{b,n}$, and ${\mathbf{\epsilon}}_{b,n}$ represent the transmit power, path loss, phase shift, and polarization pertaining to the reference signal, respectively. Notice that the LO is typically deployed close to atomic receivers. Thus, we can use line-of-sight model to generate its channel coefficients.
  • Noise model: The noise perturbing the atomic MIMO receiver is attributed to multiple sources including internal thermal noise, photon shot noise, and external interference [9]. The aggregation of sources allow us to invoke the law-of-large-number to model the noise as an additive Gaussian random vector, ${\mathbf{w}}\sim\mathcal{CN}(0,\sigma^{2}{\mathbf{I}})$.
By adding the reference signal and noise into (17), we obtain the following input-output relationship for the atomic MIMO systems as

### v3

@@ I Introduction
With the global deployment of commercial 5G services, cutting-edge research on 6G has been launched worldwide [1]. Meanwhile, the revolutionary quantum information technologies are evolving rapidly and deemed as a key enabler of the next-generation computing, communication, and sensing [2, 3, 4]. For example, ever since the invention of Schor’s algorithm, quantum computing has exhibited unparalleled abilities in quadratically or even exponentially accelerating the computing speed in solving classically intractable problems, ranging from factorization to optimization to artificial intelligence [2]. On the other hand, leveraging quantum mechanisms such as superposition and entanglement, quantum communication is capable of establishing unconditionally secure links to protect data privacy [3]. In parallel with communication and computing, quantum sensing is another promising branch of quantum information technologies [4]. It employs microscopic particles as “sensors” and capitalizes on their strong sensitivity to external disturbance, especially electromagnetic fields, to precisely measure physical quantities. Notable examples of quantum sensors include atomic clocks and superconducting interferometers. As an exciting trend, quantum sensors are being applied to an increasingly broad range of practical use cases [4, 5, 6]. Relevant techniques promise to reshape the information society and our perception of the environment. Aligned with this end, our work advocates the use of quantum sensors to revolutionize next-generation wireless receivers.
@@ I-A Concept and Properties of Atomic Receiver
Benefiting from the advancements in quantum sensing, atomic receivers offer several advantages over conventional receivers for wireless communication and sensing. First, atomic receivers can achieve much higher detection-and-sensing accuracies as mentioned earlier [4]. For example, as demonstrated in [5], atomic receivers can achieve the sub-millimeter (sub-mmWave) level granularity in detecting object vibrations using WiFi sources as opposed to the mmWave level granularity attained by traditional receivers. The second advantage lies in low thermal noise. Conventional receivers require RF circuits to intercept radio waves and downconvert RF signals to the baseband, which comprise components such as analog amplifiers and mixers that generate significant thermal noise. In contrast, atomic receivers employ all-optical components for sensing and downconversion, and thus are immune to thermal noise [9]. The above two advantages make atomic receivers a promising technology for long-range communications [6], such as satellite communications and space-air-ground networks [14], where detecting weak electric signals is crucial. Last but not the least, the flexible tunability of Rydberg atoms across an ultra-wide spectrum is a significant advantage [13]. Traditional antennas, with a fixed size typically half the wavelength of the carrier, are limited in the spectrum bandwidth they can listen to. Detecting multi-band signals requires an array of costly RF receivers, each of which is matched to a specific band [15]. In contrast, the numerous energy levels in Rydberg atoms allow an atomic receiver to respond to a broad range of frequencies from 100 Megahertz (MHz) to 1 Terahertz (THz) by exciting electrons to different energy levels [16]. This capability enables the simultaneous measurement of multi-band signals over an extremely wide spectrum using a single atomic receiver, laying the foundation for constructing an extraordinarily cost-effective full-frequency communication platform. To materialize the above visions, this work incorporates atomic receivers into modern wireless communication systems.
@@ I-B State-of-the-art Atomic Receivers
Despite the rapid development of atomic receivers in the quantum-sensing domain, their application in wireless communication is still at a nascent stage. The state-of-the-art designs are suitable only for primitive communication scenarios and schemes, such as a point-to-point system or a single-input-single-output (SISO) air interface. Without advanced designs and advices, the full potential of atomic receivers cannot be unleashed, with the state-of-the-art achieving only a Mbps-scale channel capacity [21, 16, 23]. Communication techniques that are customized for atomic receivers are largely uncharted. This calls for research on atomic-level channel coding, massive multiple-input-multiple-output (MIMO), multi-user communications, integrated sensing and communication (ISAC), reconfigurable intelligent surface (RIS), and edge learning [24, 25, 26]. To contribute to this emerging field, this paper seeks the feasibility of integrating MIMO communications and atomic receivers. MIMO, a key 5G technology, features spatial multiplexing of parallel data links and enhancement of link reliability via diversity gain [24]. It is believed that the consideration of MIMO has the potential to overcome the capacity and sensitivity bottleneck of current atomic SISO receivers.
@@ I-C Contributions and Organization
  • Performance study: We employ two metrics, namely the normalized mean square error (NMSE) and the bit error ratio (BER), to validate the feasibility of atomic MIMO receivers as well as the effectiveness of the proposed algorithms. Targeting the NMSE metric, the Cramér-Rao lower bound (CRLB) is derived. Numerical results demonstrate that the proposed EM-GS algorithm consistently approaches the CRLB and outperforms the biased GS algorithm, especially in the low signal-to-noise ratio (SNR) regime. Regarding the BER performance, simulation results confirm that the biased GS and EM-GS algorithms achieve BERs close to those from an exhaustive search, while being much more computationally efficient.
The remainder of this paper is organized as follows. The preliminaries of atomic SISO receivers are provided in Section II. The framework of atomic MIMO receiver and the corresponding detection algorithms are presented in Section III and Section IV, respectively. Section V extends the proposed designs to channel estimation problems. In Section VI, we validate the proposed design by numerical results. Finally, conclusions are drawn in Section VII.
@@ II-C1 Quantum State of Rydberg Atom
${\mathbf{E}}(t)=\bm{\epsilon}\sqrt{P}\rho s\cos(\omega t+\varphi)=[E_{x}(t),E_{y}(t),E_{z}(t)]^{T},$ (2)
@@ II-C2 State Evolution of Rydberg Atom
    $\hat{V}=\sum_{d\in\{x,y,z\}}q\hat{r}_{d}E_{d}(t)=\left(\begin{array}[]{cc}0&\xi(t)\\ \xi^{*}(t)&0\end{array}\right),$ (8)
@@
where $\delta\overset{\Delta}{=}\omega_{eg}-\omega$ denotes the detuning of the carrier frequency $\omega$ from the transition frequency $\omega_{eg}$. In (9), $\Omega_{R}\overset{\Delta}{=}\frac{|(\bm{\mu}_{eg}^{T}\bm{\epsilon})\sqrt{P}\rho s|}{\hbar}$ denotes the Rabi frequency, defined as the frequency at which the probability amplitudes of two atomic energy levels fluctuate in response to an oscillating electromagnetic field [30]. For ease of notation, we introduce $\Omega\overset{\Delta}{=}\sqrt{\Omega^{2}_{R}+\delta^{2}}$ to represent the effective Rabi frequency.
@@ II-E Implementation of Atomic Receiver
Despite the elegant mathematical representation of Rabi oscillation in (9), observing this phenomenon in practice is quite challenging due to various external disturbances that can destroy the coherence of a quantum state, such as spontaneous emission [30]. Fortunately, the data-carrying effective Rabi frequency is still measurable, as it can induce two other quantum phenomena, namely electromagnetically induced transparency (EIT) and Autler-Townes (AT) splitting [6, 8]. As shown in Fig. 2, the EIT-AT phenomenon involves a four-level system denoted by quantum states |$\ket{1}$, |$\ket{2}$, |$\ket{3}$, and |$\ket{4}$, where |$\ket{3}$ and |$\ket{4}$ represent the Rydberg states we studied before. Recall that when preparing a Rydberg state, two laser beams, the probe beam and coupling beam, are needed to excite the electron from |$\ket{1}$ to |$\ket{2}$ and then from |$\ket{2}$ to |$\ket{3}$. In practice, a PD is deployed to receive the probe beam and illustrate its spectrum for observing the Rabi frequency. When the incident electromagnetic wave is turned off and the probe beam is turned resonant with the transition |$\ket{1}\rightarrow\ket{2}$, the EIT phenomenon occurs such that the probe beam penetrates through the vapor cell nearly without absorption loss. What reflects on its spectrum is that one peak appears at the resonant frequency (see the upper sub-figure of Fig. 2). On the other hand, when the incident electromagnetic wave is turned on, the two Rydberg states |$\ket{3}$ and |$\ket{4}$ are coupled by the Rabi frequency. As a result, the AT splitting phenomenon is induced, which splits the peak of the spectrum of the probe beam into two as depicted in the lower sub-figure of Fig. 2. As physicists have demonstrated, the splitting interval $\Delta f$ of the two peaks is linearly proportional to the effective Rabi frequency [9]:
@@ III Overview of Atomic MIMO Receiver
In the preceding section, we introduced the principles of atomic SISO receiver. In this section, building on these principles, we propose the atomic MIMO receiver for multi-user communications. First, the transmission model of atomic MIMO receivers is presented. Then, the holographic-phase sensing measurement is incorporated into atomic MIMO receivers for detecting quadrature amplitude modulation (QAM) symbols, followed by the model of quantum noise.
III-A Transmission Model of Atomic MIMO Receiver
Figure 3: Schematic diagram of the atomic MIMO receiver. One atomic receiver employing $N$ atomic antennas is measuring signals from $K$-single antenna users together with one reference source.
Consider the multi-user communication system illustrated in Fig. 3, where $K$ single-antenna users communicate with one atomic MIMO receiver. Its architecture comprises $N$ vapor cells filled with Rydberg atoms and an equal number of PDs. Each vapor cell acts as an atomic antenna for measuring electromagnetic waves, while the PDs read out the measurement results (i.e., the splitting interval, $\Delta_{f}$, and the Rabi frequency, $\Omega_{R}$) and aggregate them to the digital processor for signal detection. We assume all users share the same carrier frequency $\omega$, while all atoms are triggered to the uniform Rydberg states $\hbar\omega_{g}$ and $\hbar\omega_{e}$ (e.g., $52D_{5/2}$ and $53P_{3/2}$) in response to the carrier frequency (e.g., $\omega=5$ GHz). As a result, the resonance case with $\omega=\omega_{eg}$ is created.
Consider an arbitrary user, say the $k$-th user. Its phase modulated symbol is denoted as $s_{k}=|s_{k}|e^{i\gamma_{k}}\in\mathcal{S}$, where the set $\mathcal{S}$ corresponds to the QAM constellation. The average power of $s_{k}$ is normalized as $\mathsf{E}(\|s_{k}\|^{2})=1$. Then, the transmit RF signal is given by $x_{k}(t)=\sqrt{P_{k}}|s_{k}|\cos(\omega t+\gamma_{k})$. We adopt the practical multi-path wireless channel with $L$ paths. The simultaneous transmissions from all users result in the received electromagnetic wave at the $n$-th vapor cell:
@@ III-A Transmission Model of Atomic MIMO Receiver
$\color[rgb]{0,0,0}\xi_{n}(t){=}\sum_{k=1}^{K}\sum_{l=1}^{L}{\bm{\mu}}_{eg}^{T}\bm{\epsilon}_{nkl}\sqrt{P_{k}}\rho_{nkl}|s_{k}|\cos(\omega t+\varphi_{nkl}+\gamma_{k}).$ (14)
@@
$\color[rgb]{0,0,0}\Omega_{R,n}=\left|\sum_{k=1}^{K}\sum_{l=1}^{L}\frac{1}{\hbar}({\bm{\mu}}_{eg}^{H}\bm{\epsilon}_{nkl})\sqrt{P_{k}}\rho_{nkl}|s_{k}|e^{i(\varphi_{nkl}+\gamma_{k})}\right|.$ (16)
@@
The goal of atomic-MIMO receiver is to read the values of $\{\Omega_{R,n}\}$ and infer $\{s_{k}\}$ from $\{\Omega_{R,n}\}$. To this end, according to the principle of EIT-AT discussed in Section II-E, the Rabi frequency, $\Omega_{R,n}$, can be measured by the splitting interval, $\Delta f_{n}$, of the probe-beam spectrum since $\Omega_{R,n}=2\pi\frac{\lambda_{p}}{\lambda_{c}}\Delta f_{n}$. For ease of exposition, let the measurement of $\Omega_{R,n}$ be denoted as $z_{n}$, which is the equivalent received signal at the $n$-th antenna. To facilitate the detector design, a compact matrix form of $\{z_{n}\}$ is derived as follows. We define ${\mathbf{s}}=[s_{1},s_{2},\cdots,s_{K}]^{T}$ and ${\mathbf{a}}_{n}=[a_{n,1},a_{n,2},\cdots,a_{n,K}]^{T}$, where $a_{n,k}=\sum_{l=1}^{L}\frac{1}{\hbar}({\bm{\mu}}_{eg}^{T}\bm{\epsilon}_{n,k})\sqrt{P_{k}}\rho_{nkl}e^{-i\varphi_{nkl}}$. It is straightforward to obtain that $z_{n}=\left|{\mathbf{a}}_{n}^{H}{\mathbf{s}}\right|$. We further define the received vector-form signal as ${\mathbf{z}}=[z_{1},z_{2},\cdots,z_{N}]^{T}$ and the observation matrix as ${\mathbf{A}}=[{\mathbf{a}}_{1},{\mathbf{a}}_{2},\cdots,{\mathbf{a}}_{N}]\in\mathbb{C}^{K\times N}$. It follows that
@@
where the operator $|{\mathbf{x}}|$ takes the magnitude of each element of the vector, ${\mathbf{x}}$. A crucial conclusion is drawn from (17) that the transmit symbols in ${\mathbf{s}}$ are embedded in the observed amplitude vector ${\mathbf{z}}=|{\mathbf{A}}^{H}{\mathbf{s}}|$, where the observation matrix ${\mathbf{A}}$ can be treated as an effective MIMO channel. As opposed to the linear model adopted in conventional MIMO communications, the atomic MIMO receiver features a non-linear model based on amplitude detection. This is equivalent to the well-known PR model in the realm of optical imaging [31]. Before delving into this PR problem, two factors, the holographic-phase sensing measurement and the noise model, that are part of the proposed framework and affect the receiver’s effectiveness should be discussed.
III-B Holographic-Phase Sensing Measurement
One can observe from (17) that the absolute phase $e^{i\angle{\mathbf{s}}}$ of ${\mathbf{s}}$ is ambiguous to the received signal ${\mathbf{z}}$, making it hard to detect QAM symbols. To address this issue, a measurement scheme called holographic phase-sensing is often employed [20]. As illustrated in Fig. 3, this scheme artificially introduces a LO to send known reference signals, denoted as $s_{b}=|s_{b}|e^{j\gamma_{b}}$, to interfere with signals from users. The LO is typically implemented by a horn antenna and can be deployed close to the receiver. Utilizing the reference signal $s_{b}$, one can infer the absolute phase, $e^{i\angle{\mathbf{s}}}$, by estimating the phase difference between ${\mathbf{s}}$ and $s_{b}$, thereby enabling QAM detection.
Mathematically speaking, let the reference signal share the same carrier frequency $\omega$ with the user signal. The transmit RF signal by LO is expressed as $x_{b}(t)=\sqrt{P_{b}}|s_{b}|\cos(\omega t+\gamma_{b})$. Since the LO-to-receiver distance is quite short, e.g., tens of centimeters, the LO-to-receiver channel is dominated by the line-of-sight (LoS) path. Thereby, the received electromagnetic wave in (11) should be modified as
${\mathbf{E}}_{n}(t)=\sum_{k=1}^{K}\sum_{l=1}^{L}\bm{\epsilon}_{nkl}\sqrt{P_{k}}\rho_{nkl}|s_{k}|\cos(\omega t+\varphi_{nkl}+\gamma_{k})$
$+\bm{\epsilon}_{b,n}\sqrt{P_{b}}\rho_{b,n}|s_{b}|\cos(\omega t+\varphi_{b,n}+\gamma_{b}),$ (18)
where $\rho_{b,n}$, $\varphi_{b,n}$, and $\bm{\epsilon}_{b,n}$ refer to the path loss, phase shift, and polarization of the LO-to-receiver LoS path, respectively. Then, following the same steps as (12) to (17), we can arrive at the modified received signal:
${\mathbf{z}}=|{\mathbf{A}}^{H}{\mathbf{s}}+{\mathbf{b}}|,$ (19)
where ${\mathbf{b}}=[b_{1},\cdots,b_{N}]^{T}$ and $b_{n}=\frac{s_{b}}{\hbar}{\bm{\mu}}_{eg}^{H}\bm{\epsilon}_{b,n}\sqrt{P_{b}}\rho_{b,n}e^{i\varphi_{b,n}}$. Due to the presence of reference signal ${\mathbf{b}}$, (19) is called a biased PR model as opposed to the typical PR model without this reference signal.
III-C Noise Model
The noise perturbing the atomic MIMO receiver is attributed to both external and internal sources.
External noise is mainly induced by the signal interference from adjacent time slots, bands, and nearby cells, which is clearly environment-dependent. We use the notation $\sigma_{E}^{2}$ to represent the total variance of external noise.
The internal noise comes from the thermal noise of electronic components and the randomness of quantum measurement, i.e., quantum shot noise (QSN). In contrast to conventional RF systems, an atomic receiver makes no use of RF circuits at the front end that produce thermal noises, such as analog filters, amplifiers, and mixers. This fact means that an atomic receiver is immune to thermal noise. Thereafter, the limiting noise source for atomic receivers is the QSN, whose level in the unit of electric field strength [V/m] is presented by [6]
$E_{\min}=\frac{h}{\|\bm{\mu}_{eg}\|(T_{i}T_{r}N_{a})^{1/2}},$ (20)
where $T_{i}$, $T_{r}$, and $N_{a}$ denote the measurement time, the relaxation time of Rydberg states, and the number of participating atoms, respectively. By plugging this noise limit into the electric field and further transforming it into the unit of Rabi frequency [Rad/s], $E_{\min}$ is normalized to $\frac{\|\bm{\mu}_{eg}\|}{\hbar}E_{\min}=\frac{2\pi}{(T_{t}T_{r}N_{a})^{1/2}}$. As a result, the variance of QSN is given as
$\sigma_{S}^{2}=\frac{4\pi^{2}}{T_{i}T_{r}N_{a}}.$ (21)
The extreme sensitivity of atomic receivers arises because the noise level (21) can be orders of magnitude lower than the thermal noise of conventional receivers [5]. To fairly compare the quantum shot noise and thermal noise, we adopt the standard $T_{i}=1\>{\rm s}$ measurement time for both them. In this case, the power density of room-temperature thermal noise is $-176\>{\rm dBm}$ [32]. As for the quantum noise, we adopt the typical parameters $T_{r}=10\>\mu{\rm s}$ and $N_{a}=5\times 10^{5}$, which give rise to a noise level $\frac{2\pi}{(T_{i}T_{r}N_{a})^{1/2}}=2.81\>{\rm Hz}$. According to transformation between electric field and power [32], this noise refers to a power density of $-191\>{\rm dBm}$ when $\omega=2\pi\times 5\>{\rm GHz}$, which is 15 dB smaller than the room-temperature thermal noise level. This reflects the high sensitivity of atomic receivers [5].
In summary, the total variance of noise is expressed as
$\sigma^{2}=\sigma_{S}^{2}+\sigma_{E}^{2}=\frac{4\pi^{2}}{T_{i}T_{r}N_{a}}+\sigma_{E}^{2}.$ (22)
We invoke the law-of-large-number to model the noise as an additive Gaussian random vector, ${\mathbf{w}}\sim\mathcal{CN}(0,\sigma^{2}{\mathbf{I}})$. By adding the noise into (III-B), we finally arrive at the following input-output relationship for atomic MIMO systems: